OpenStudy (anonymous):

What is the common ratio of this geometric sequence? 1.1, 4.4, 17.6, 70.4,... A. 4 B. 11 C. 64

OpenStudy (anonymous):

A = 4 common ration = a(n+1) / a(n) = 4.4 / 1.1 = 17.6 / 4.4 = 4

OpenStudy (anonymous):

4.4 / 1.1 = 17.6 /4.4 = 70.4/17.6 = 4 common ratio is 4